# GTR: Guided Thought Reinforcement Prevents Thought Collapse in RL-based VLM Agent Training

### v1

@@ Abstract
Figure 1: Zoom in for more details. Illustration of thought collapse occurring when training a VLM agent to solve 24 points through RL. With RL training, the thoughts generated by the VLM agent quickly lose their diversity, turn out to be incorrect, and lead to invalid actions and negative rewards. For example, checkpoints ➌ and ➍ erroneously predict the same thought (“I should append ‘10’ to the current formula”) and action 10 even in the face of different environment states, catastrophically degrading the agent’s decision-making capabilities and RL’s sample efficiency. We propose Guided Thought Reinforcement (GTR) to prevent this problem.
@@ 1 Introduction
The rapid evolution of large language models (LLMs) and vision-language models (VLMs) has significantly enhanced machines’ ability to comprehend general text and images. Through a comprehensive reinforcement learning (RL) training pipeline, these models can be improved based on evaluations (i.e., rewards) of their outcomes, thus aligning with human values, emerging reasoning capabilities via long chain-of-thought (CoT), and achieving higher success rates across various tasks. Yet, few studies have explored how to train VLM agents in dynamic visual environments to infer a sequence of correct actions based on their perceived information and ultimately accomplish specific goals.
Although prior work [61] has demonstrated the feasibility of RL with verifiable outcome rewards (RLVR) for fine-tuning VLM agents to solve multi-step decision-making tasks, their progress remains limited in environments involving longer episodes, larger state spaces, and more substantial reasoning requirements. In experiments carried out on representative tasks such as the 24 points card game [61] and the embodied environment ALFWorld [38], we identified the bottleneck that restricts further emergence of the model’s decision-making and reasoning capabilities, which we refer to as thought collapse. In RL training, rewards are entirely derived from the final action, without considering the intermediate reasoning thoughts (see Fig. 1 for more details). When facing challenging and complex tasks, the guiding effect of action rewards is primarily weakened. The agent’s CoT process rapidly loses diversity, leading to incorrect, incoherent, and rigid thoughts even with different visual and textual inputs, ultimately leading to erroneous actions and negative rewards. Although the agent continues to generate reasoning thoughts, it has already lost the ability to think, hindering the emergence of its full potential, as illustrated by the checkpoints ➌ and ➍ in Fig. 1.
@@
Specifically, as shown in Fig. 3, we design a plug-and-play VLM corrector model built upon any off-the-shelf VLM to evaluate and refine the agent’s thoughts at each RL training step, thus automating the correction of collapsed thought trajectories. Inspired by previous works that integrate guidance into RL training [18, 9, 35], our GTR framework allows the agent to perform SFT thought-cloning alongside PPO updates, ensuring both the rationality of the reasoning process and the correctness of the final actions. Furthermore, we address the issue of output format degradation with format rewards and repetition penalties, enhance the accuracy and coherence of thought correction through appropriate tool usage, and mitigate distribution shift in the thought-cloning process by a prevalent imitation learning method Dataset Aggregation (DAgger) [33]. The integration of thinking process guidance allows the VLM agent to generate a structured and reliable reasoning and action trajectory, leading to more transparent and interpretable decision-making in complex and challenging visual environments.
Empirically, we apply GTR to train an LLaVA-7b model [22, 21, 23]. On the highly challenging card games such as 24 points, GTR achieved over 300% task success rate and significantly higher returns compared to the state-of-the-art (SoTA) methods (including the agents driven by Qwen-VL-2.5, Gemini, and GPT-4o), demonstrating that combining thought guidance with RL unleashes the decision-making potential of VLM agents more effectively. Additionally, in experiments on the embodied environment ALFWorld, GTR exhibits a better success rate and sample efficiency, proving the generality of our framework across diverse tasks.
@@ 2.1 LLMs/VLMs as Decision-making Agents
Recent advancement in foundation models has expanded their applications beyond text and image generation to complex decision-making tasks, focusing on their ability to reason, plan, and interact with dynamic environments [51, 55, 19, 56]. Many previous works have leveraged prompting techniques to activate decision-making capabilities from frozen LLMs [49, 57, 58, 48, 46, 16, 29, 2, 37, 50]. Furthermore, agents with VLMs as backbones can perform reasoning based on visual inputs, either by translating their observations to text descriptions or aligning their embeddings with LLMs [2, 7, 11, 17, 25, 40, 54, 3, 63]. However, these approaches either do not involve model adjustments or are limited to pretraining additional layers on fixed datasets, making them unable to adapt to dynamic environments through interaction. In contrast, we explore reinforcement learning to finetune the entire VLM, enhancing the model’s ability to reason and make decisions by interacting with the environment through open-ended text.
@@ 3 Thought Collapse
Our experiments may identify the root of this problem: the agent’s thought process often becomes irrational or templated during RL training, significantly impairing its decision-making abilities. As illustrated in Fig. 1, the agent fails to improve task success rates or episode returns. Instead, it generates fragmented thoughts and degenerates into a strategy of producing similar outputs for different states. Although the agent continues to output thoughts and action decisions at this stage, it has clearly lost its ability to engage in meaningful reasoning and decision-making. We term this catastrophic phenomenon as “thought collapse”.
@@ 4.1 RL Training of VLM Agents
The action probability required for policy gradient is calculated from the generation probabilities of each token in the output text. Additionally, a scaling factor is employed to balance the longer length of CoT outputs compared to action outputs. If $\pi_{\theta}$ denotes the policy, $s_{t}$ and $a_{t}$ represent the state and action, $\boldsymbol{v}_{t}^{\text{in}}$ as input tokens, $\boldsymbol{v}_{t}^{\text{tht}}$ and $\boldsymbol{v}_{t}^{\text{act}}$ represent CoT reasoning tokens and action tokens respectively, this calculation is shown as:
$\log\pi_{\theta}(a_{t}|s_{t},\boldsymbol{v}_{t}^{\text{in}})$ (2)
$=$ $\lambda\log\pi_{\theta}(\boldsymbol{v}_{t}^{\text{tht}}|o_{t},\boldsymbol{v}_{t}^{\text{in}})+\log\pi_{\theta}(\boldsymbol{v}_{t}^{\text{act}}|o_{t},\boldsymbol{v}_{t}^{\text{in}})$
$=$ $\lambda\sum\log\frac{p(o_{t},\boldsymbol{v}_{t}^{\text{in}},\boldsymbol{v}^{\text{tht}}_{[:i]})}{p(o_{t},\boldsymbol{v}_{t}^{\text{in}},\boldsymbol{v}^{\text{tht}}_{[:i-1]})}+\sum\log\frac{p(o_{t},\boldsymbol{v}_{t}^{\text{in}},\boldsymbol{v}^{\text{act}}_{[:i]})}{p(o_{t},\boldsymbol{v}_{t}^{\text{in}},\boldsymbol{v}^{\text{act}}_{[:i-1]})}.$
@@ 4.2 Process Guidance from a VLM Corrector
Figure 4: Comparison among different process guidance methods in the 24 points card game. Trivial numerical rewards provided by the VLM judge and rule-based evaluation cannot incentivize the agent’s reasoning thoughts and higher success rate.
Various approaches that introduce process guidance to model training have been explored. Process Reward Models (PRMs) represent a traditional and widely used method for process supervision, particularly in mathematical reasoning tasks with LLMs [44, 20]. However, high-quality vision-language data annotation is extremely expensive and time-consuming. Moreover, learning with static datasets cannot cover a vast variety of decision-making scenarios, often leading to biased policies when deployed to dynamic environments. An intuitive solution to the above challenges is the VLM-as-a-judge approach, which induces a VLM to score the agent’s outputs [62, 10, 52]. However, this method does not demonstrate reasonable performance (see Fig. 4). While episodic returns increase to some extent, task success rates fail to improve. This is because using naive numerical scores does not provide sufficient and accurate guidance for effective RL training, especially given the strong reward-hacking capabilities of large models. In challenging tasks where the model’s baseline ability is weak, the lack of positive incentives can cause the agent to fall into passive exploration, further hindering performance. Similar issues also arise with methods like length-based rewards [59], which aim to encourage longer thoughts. We find that the guidance provided by these approaches is not enough for the emergence of reasoning capabilities.
These limitations call for a simple, scalable, and informative process guidance framework, and we demonstrate the corrector model as a promising solution. Our GTR framework leverages an external VLM as a corrector model, which first evaluates the agent’s thought at each step for visual recognition accuracy and reasoning correctness, then refines the original thought. When the corrector identifies inconsistencies or errors, the model performs corrections based on the agent’s outputs. To incorporate corrected thoughts into the agent while preserving scalability, we draw inspiration from the recent work [15] and add a simple SFT loss over the thought tokens $\boldsymbol{v}_{t}^{\text{tht}}$ to the PPO loss, aligning the agent’s reasoning with the corrected thought trajectories, as shown in Fig 3. Formally, if we denote the agent’s thought output as $th$ and action as $a$, given agent model $\pi_{\theta}$ and corrector model $\pi_{\text{corr}}$, the objective of GTR can be represented as:
$\min_{\theta}\mathop{\mathbb{E}}_{{s,(th,a)}\sim\pi_{\theta}}\left[\mathcal{L}_{\text{PPO}}(s,a)+\mathcal{L}_{\text{SFT}}(s,\pi_{\text{corr}}(s,th))\right],$ (3)
@@
$\mathcal{L}_{\text{PPO}}(s,a)$ $=-\min\left(\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{k}}(a|s)}A^{\pi_{\theta_{k}}}(s,a),\right.$
$\left.\text{clip}\left({\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{k}}(a|s)},1+c,1-c}\right)A^{\pi_{\theta_{k}}}(s,a)\right),$
$\mathcal{L}_{\text{SFT}}(s,th)$ $=-\sum_{t}\log P(th_{t}|s,th_{<t};\theta).$ (4)
@@ 4.3 Mitigate Distribution Shift in Thought Cloning
Research such as TD3+BC [9], which also integrates supervised signals with reinforcement learning, typically focuses on the off-policy or offline RL algorithms. However, we encountered a distribution shift issue when incorporating thought cloning into the PPO training of VLM agents. As the agent’s policy iteratively updates, the PPO algorithm discards previous data and resamples in each round. Performing thought cloning on this non-i.i.d dataset can lead to catastrophic forgetting or error accumulation. To address this, we adopt an interactive imitation learning algorithm, Dataset Aggregation (DAgger) [33], aggregating all historical corrections and samples from them. This approach has been proven to converge to the expert policy, in this case, the outputs of the corrector model. If $\mathcal{B}$ represents the PPO data buffer and $\mathcal{D}$ denotes the DAgger thought dataset, the Eq. 3 can be rewritten as below.
@@
8:         $\boldsymbol{v}_{t}^{\text{in}}=h(s_{t})$triangleright Obtain prompt
@@
14:         $\mathcal{B}\leftarrow\mathcal{B}\cup(s_{t},a_{t},r_{t},\boldsymbol{v}_{t}^{\text{out}},\log\pi_{\theta}(a_{t}|o_{t},\boldsymbol{v}_{t}^{\text{in}}))$
15:         $\mathcal{D}\leftarrow\mathcal{D}\cup(s_{t},\boldsymbol{\hat{v}}_{t}^{\text{tht}})$
@@ 4.4 Improving Data Quality
We observe that RL training, which requires the model to explore and discover better outputs, also introduces the risk of generating poor outputs. During the RL finetuning of VLMs, it is common for the model to produce an invalid format or exhibit excessive repetition, which has also been noted in other recent studies [12, 13]. Therefore, we incorporate a repetition penalty during the model’s response generation process and explicitly integrate format judgment into the corrector model so that answers with a valid format receive a format reward at each step. These measures significantly improve the stability of the model’s output format.
While general-purpose VLM correction models possess extensive general knowledge and reasoning capabilities, they may lack task-specific expertise. To this end, we leverage the function-calling ability of large models, enabling the corrector model to access specific information or function modules during the evaluation and correction process. This further enhances the accuracy and credibility of corrections. For instance, in the 24 points game, the corrector may invoke a piece of Python code that calculates possible equations that evaluate to 24, given the model’s identification of card ranks in the picture. The introduction of corrector models also enables data sampling control within our framework. By referencing the corrector’s judgments, we can truncate episodes when the agent enters unsolvable or meaningless states, thereby improving the data quality in the buffer and assisting RL training efficiency. In the 24-point game, where each action is irreversible, the episode is ended if the four cards cannot accomplish 24 or if the agent’s historical outputs eliminate the possibility of finishing the task. In ALFWorld, we apply truncation when the agent’s action sequence becomes excessively long or when it continuously repeats the same ineffective action.
@@ 5.1 Experimental Setup
We mainly base our experiments on the Points24 task within the $\mathtt{gym\_cards}$ environment [61], which requires the model to perform fine-grained visual recognition of poker cards and then engage in language reasoning. Tasks other than Points24 require fewer recognitions, significantly smaller state and action spaces, and shorter action sequences (typically solvable within five steps). As a result, outcome rewards are sufficient to guide RL training and thought collapse rarely occurs. In contrast, the correct action sequence of Points24 can exceed 10 steps, and the combination of a long formula and larger action space creates a much broader decision-making space. Previous methods have not achieved satisfactory results. Therefore, our experiments focus primarily on this complex problem.
@@
To validate the generality of thought collapse and the GTR framework, we also evaluate performance on the ALFWorld benchmark, a multimodal simulation platform featuring various embodied household tasks [38]. These tasks are divided into six categories: Pick & Place, Clean & Place, Heat & Place, Cool & Place, Look in Light, and Pick Two Objects & Place. Given the current visual observation, the agent needs to determine the following action following a predefined instruction, such as “go to cabinet 1”, “open drawer 1” or “cool apple 1 with fridge 1”. The tasks involve navigation and interaction with objects in the environment, presenting significant challenges in visual recognition, long-term planning, and commonsense reasoning.
Notably, ALFWorld provides visual and text observations of the environment, and previous works have used both as input, reducing the difficulty of visual recognition. We observe with various models that this setting often leads the agent to rely heavily on text descriptions while neglecting visual input. Therefore, to evaluate the comprehensive multimodal decision-making capability of agents and to better simulate real-world scenarios, we removed the text environment description in our experiments. We also incorporate historical actions as part of the input, making it more suitable to assess the sequential decision-making capabilities of agents.
@@
Our primary baseline is RL4VLM [61], which directly uses environmental rewards for PPO training. It is the first framework to finetune VLM agents with RL and represents the SoTA method. Accordingly, we consider the method that performs only thought cloning, without RL (SFT-only). When evaluating final performance, we also include the LLaVA-mistral-7b base model and commercial API-based models. For Alfworld, the challenging environment with navigation and household tasks, we collect an expert dataset using script policy and fine-tune several SoTA VLM models based on this dataset, using them as baselines.
@@
The GTR framework leverages a corrector model with established capabilities for automating thought correction. In our experiment, we select the GPT-4o model for this role. Consistent with previous research, to ensure that RL training begins with a model possessing reasonable instruction following abilities, all training starts from an SFT-initialized LLaVA-mistral-7b model. We train the model for 15k steps on the Points24 task and 5k steps on the ALFWorld task, aligning with baseline settings. On a single A100 (40GB) GPU, the training process with LoRA [14] takes approximately 30 hours to complete.

### v2

@@ 1 Introduction
The rapid evolution of large language models (LLMs) and vision-language models (VLMs) has significantly enhanced machines’ ability to comprehend general text and images. Through a comprehensive reinforcement learning (RL) training pipeline, these models can be improved based on evaluations (i.e., rewards) of their outcomes, thereby aligning with human values, emerging reasoning capabilities via long-chain-of-thought (CoT), and achieving higher success rates across various tasks. Yet, few studies have explored how to train VLM agents in dynamic visual environments to infer a sequence of correct actions based on their perceived information and ultimately accomplish specific goals.
Although prior work [61] has demonstrated the feasibility of RL with verifiable outcome rewards (RLVR) for fine-tuning VLM agents to solve multi-step decision-making tasks, their progress remains limited in environments involving longer episodes, larger state spaces, and more substantial reasoning requirements. In experiments conducted on representative tasks, such as the 24 points card game [61] and the embodied environment ALFWorld [38], we identified the bottleneck that restricts the further emergence of the model’s decision-making and reasoning capabilities, which we refer to as thought collapse. In RL training, rewards are entirely derived from the final action, without considering the intermediate reasoning thoughts (see Fig. LABEL:fig:thought_collapse for more details). When facing challenging and complex tasks, the guiding effect of action rewards is primarily weakened. The agent’s CoT process rapidly loses diversity, resulting in incorrect, incoherent, and rigid thoughts, even with different visual and textual inputs, ultimately leading to erroneous actions and negative rewards. Although the agent continues to generate reasoning thoughts, it has already lost the ability to think, hindering the emergence of its full potential, as illustrated by the checkpoints ➌ and ➍ in Fig. LABEL:fig:thought_collapse.
@@
Specifically, as shown in Fig. 2, we design a plug-and-play VLM corrector model built upon any off-the-shelf VLM to evaluate and refine the agent’s thoughts at each RL training step, thus automating the correction of collapsed thought trajectories. Inspired by previous works that integrate guidance into RL training [18, 9, 35], our GTR framework allows the agent to perform SFT thought-cloning alongside PPO updates, ensuring both the rationality of the reasoning process and the correctness of the final actions. Furthermore, we address the issue of output format degradation with format rewards and repetition penalties, enhance the accuracy and coherence of thought correction through appropriate tool usage, and mitigate distribution shift in the thought-cloning process by a prevalent imitation learning method, Dataset Aggregation (DAgger) [33]. The integration of thinking process guidance enables the VLM agent to generate a structured and reliable reasoning and action trajectory, resulting in more transparent and interpretable decision-making in complex and challenging visual environments.
Empirically, we apply GTR to train an LLaVA-7B model [22, 21, 23]. On the highly challenging card games such as 24 points, GTR achieved over 300% task success rate and significantly higher returns compared to the state-of-the-art (SoTA) methods (including the agents driven by Qwen2-VL, Gemini, and GPT-4o), demonstrating that combining thought guidance with RL unleashes the decision-making potential of VLM agents more effectively. Additionally, in experiments on the embodied environment ALFWorld, GTR exhibits a better success rate and sample efficiency, proving the generality of our framework across diverse tasks.
@@ 2.1 LLMs/VLMs as Decision-making Agents
Recent advancement in foundation models has expanded their applications beyond text and image generation to complex decision-making tasks, focusing on their ability to reason, plan, and interact with dynamic environments [51, 55, 19, 56]. Many previous works have leveraged prompting techniques to activate decision-making capabilities from frozen LLMs [49, 57, 58, 48, 46, 16, 29, 2, 37, 50]. Furthermore, agents with VLMs as backbones can perform reasoning based on visual inputs, either by translating their observations to text descriptions or aligning their embeddings with LLMs [2, 7, 11, 17, 25, 40, 54, 3, 63]. However, these approaches either do not involve model adjustments or are limited to pretraining additional layers on fixed datasets, making them unable to adapt to dynamic environments through interaction. In contrast, we explore reinforcement learning to fine-tune the entire VLM, enhancing the model’s ability to reason and make decisions by interacting with the environment through open-ended text.
@@ 3 Thought Collapse
Our experiments may identify the root of this problem: the agent’s thought process often becomes irrational or templated during RL training, significantly impairing its decision-making abilities. As illustrated in Fig. LABEL:fig:thought_collapse, the agent fails to improve task success rates or episode returns. Instead, it generates fragmented thoughts and degenerates into a strategy of producing similar outputs for different states. Although the agent continues to output thoughts and action decisions at this stage, it has clearly lost its ability to engage in meaningful reasoning and decision-making. We term this catastrophic phenomenon as “thought collapse”.
@@ 4.1 RL Training of VLM Agents
[…]
$\log\pi_{\theta}(a_{t}|o_{t},\boldsymbol{v}_{t}^{\text{in}})$ (2)
$=$ $\lambda\log\pi_{\theta}(\boldsymbol{v}_{t}^{\text{tht}}|o_{t},\boldsymbol{v}_{t}^{\text{in}})+\log\pi_{\theta}(\boldsymbol{v}_{t}^{\text{act}}|o_{t},\boldsymbol{v}_{t}^{\text{in}},\boldsymbol{v}_{t}^{\text{tht}})$
$=$ $\lambda\sum\log\frac{p(o_{t},\boldsymbol{v}_{t}^{\text{in}},\boldsymbol{v}^{\text{tht}}_{[:i]})}{p(o_{t},\boldsymbol{v}_{t}^{\text{in}},\boldsymbol{v}^{\text{tht}}_{[:i-1]})}+\sum\log\frac{p(o_{t},\boldsymbol{v}_{t}^{\text{in}},\boldsymbol{v}_{t}^{\text{tht}},\boldsymbol{v}^{\text{act}}_{[:i]})}{p(o_{t},\boldsymbol{v}_{t}^{\text{in}},\boldsymbol{v}_{t}^{\text{tht}},\boldsymbol{v}^{\text{act}}_{[:i-1]})}.$
@@ 4.2 Process Guidance from a VLM Corrector
Various approaches that introduce process guidance to model training have been explored. Process Reward Models (PRMs) represent a traditional and widely used method for process supervision, particularly in mathematical reasoning tasks with LLMs [44, 20]. However, high-quality vision-language data annotation is extremely expensive and time-consuming. Moreover, learning with static datasets cannot cover a vast variety of decision-making scenarios, often leading to biased policies when deployed to dynamic environments. An intuitive solution to the above challenges is the VLM-as-a-judge approach, which induces a VLM to score the agent’s outputs [62, 10, 52]. However, this method does not demonstrate reasonable performance (see Fig. 3). While episodic returns increase to some extent, task success rates fail to improve. This is because using naive numerical scores does not provide sufficient and accurate guidance for effective RL training, especially given the strong reward-hacking capabilities of large models. In challenging tasks where the model’s baseline ability is weak, the lack of positive incentives can cause the agent to fall into passive exploration, further hindering performance. Similar issues also arise with methods like length-based rewards [59], which aim to encourage longer thoughts.
Figure 3: Comparison among different process guidance methods in the 24 points card game. Trivial numerical rewards provided by the VLM judge or rule-based evaluation cannot incentivize the agent’s reasoning thoughts and higher success rate.
These limitations call for a simple, scalable, and informative process guidance framework, and we demonstrate the corrector model as a promising solution. Our GTR framework leverages an external VLM as a corrector model, which first evaluates the agent’s thought at each step for visual recognition accuracy and reasoning correctness, then refines the original thought. When the corrector identifies inconsistencies or errors, the model performs corrections based on the agent’s outputs. To incorporate corrected thoughts into the agent while preserving scalability, we draw inspiration from the recent work [15, 56] and add a simple SFT loss over the thought tokens $\boldsymbol{v}_{t}^{\text{tht}}$ to the PPO loss, aligning the agent’s reasoning with the corrected thought trajectories, as shown in Fig. 2. Formally, if we denote the agent’s thought output as $th$ and action as $a$, given agent model $\pi_{\theta}$ and corrector model $\pi_{\text{corr}}$, the objective of GTR can be represented as:
$\min_{\theta}\mathop{\mathbb{E}}_{{o,(th,a)}\sim\pi_{\theta}}\left[\mathcal{L}_{\text{PPO}}(o,a)+\mathcal{L}_{\text{SFT}}(o,\pi_{\text{corr}}(o,th))\right],$ (3)
@@
$\mathcal{L}_{\text{PPO}}(o,a)$ $=-\min\left(\frac{\pi_{\theta}(a|o)}{\pi_{\theta_{k}}(a|o)}A^{\pi_{\theta_{k}}}(o,a),\right.$
$\left.\text{clip}\left({\frac{\pi_{\theta}(o|s)}{\pi_{\theta_{k}}(o|s)},1+c,1-c}\right)A^{\pi_{\theta_{k}}}(o,a)\right),$
$\mathcal{L}_{\text{SFT}}(o,th)$ $=-\sum_{t}\log P(th_{t}|o,th_{<t};\theta).$ (4)
@@ 4.3 Mitigate Distribution Shift in Thought Cloning
Research such as TD3+BC [9], which also integrates supervised signals with reinforcement learning, typically focuses on the off-policy or offline RL algorithms. However, we encountered a distribution shift issue when incorporating thought cloning into the PPO training of VLM agents. As the agent’s policy iteratively updates, the PPO algorithm discards previous data and resamples in each round. Performing thought cloning on this non-i.i.d dataset can lead to catastrophic forgetting or error accumulation. To address this, we adopt an interactive imitation learning algorithm, Dataset Aggregation (DAgger) [33], aggregating all historical corrections and samples from them. This approach has been proven to converge to the expert policy, specifically the outputs of the corrector model. If $\mathcal{B}$ represents the PPO data buffer and $\mathcal{D}$ denotes the DAgger thought dataset, Eqn. 3 can be rewritten as below.
@@
8:         $\boldsymbol{v}_{t}^{\text{in}}=h(o_{t})$triangleright Obtain prompt
@@
14:         $\mathcal{B}\leftarrow\mathcal{B}\cup(o_{t},a_{t},r_{t},\boldsymbol{v}_{t}^{\text{out}},\log\pi_{\theta}(a_{t}|o_{t},\boldsymbol{v}_{t}^{\text{in}}))$
15:         $\mathcal{D}\leftarrow\mathcal{D}\cup(o_{t},\boldsymbol{\hat{v}}_{t}^{\text{tht}})$
@@ 4.4 Improving Data Quality
We observe that RL training, which requires the model to explore and discover better outputs, also introduces the risk of generating poor outputs. During the RL fine-tuning of VLMs, it is common for the model to produce an invalid format or exhibit excessive repetition, a phenomenon also noted in other recent studies [12, 13]. Therefore, we incorporate a token-level repetition penalty during the model’s response generation process and explicitly integrate format judgment into the corrector model so that answers with a valid format receive a format reward at each step. These measures significantly improve the stability of the model’s output format.
While general-purpose VLM correction models possess extensive general knowledge and reasoning capabilities, they may lack task-specific expertise. To this end, we leverage the function-calling ability of large models, enabling the corrector model to access specific information or function modules during the correction process. This further enhances the accuracy and credibility of corrections. For instance, in the 24-point game, the corrector may invoke a piece of Python code that calculates possible equations that evaluate to 24. The introduction of corrector models also enables data sampling control. By referencing the corrector’s judgments, we can truncate episodes when the agent enters unsolvable or meaningless states, thereby improving the data quality in the buffer and assisting RL training efficiency. In the 24-point game, the episode ends if the agent is impossible to complete the task. In ALFWorld, we apply truncation when the agent’s action sequence becomes excessively long or when it continuously repeats the same ineffective action.
@@ 5.1 Experimental Setup
We mainly base our experiments on the Points24 task within the $\mathtt{gym\_cards}$ environment [61], which requires the model to perform fine-grained visual recognition of poker cards and then engage in language reasoning. Tasks other than Points24 require fewer recognitions, significantly smaller state and action spaces, and shorter action sequences (typically solvable within five steps). As a result, outcome rewards are sufficient to guide RL training, and thought collapse rarely occurs. In contrast, the correct action sequence of Points24 can exceed 10 steps, and the combination of a long formula and larger action space creates a much broader decision-making space. Previous methods have not achieved satisfactory results. Therefore, our experiments focus primarily on this complex problem.
@@
To validate the generality of thought collapse and the GTR framework, we also evaluate performance on the ALFWorld benchmark, a multimodal simulation platform featuring various embodied household tasks [38]. These tasks are divided into six categories: Pick & Place, Clean & Place, Heat & Place, Cool & Place, Look in Light, and Pick Two Objects & Place. Given the current visual observation, the agent needs to determine the next action following a predefined instruction, such as “go to cabinet 1”, “open drawer 1” or “cool apple 1 with fridge 1”. The tasks involve navigation and interaction with objects in the environment, presenting significant challenges in visual recognition, long-term planning, and commonsense reasoning.
Notably, ALFWorld provides visual and textual observations of the environment, and previous works have utilized both as input, thereby reducing the difficulty of visual recognition. We observe with various models that this setting often leads the agent to rely heavily on text descriptions while neglecting visual input. Therefore, to evaluate the comprehensive multimodal capability of agents and to better simulate real-world scenarios, we removed the text description in our experiments. We also incorporate historical actions as part of the input, making it more suitable to assess the sequential decision-making capabilities of agents.
@@
Our primary baseline is RL4VLM [61], which directly uses environmental rewards for PPO training. It is the first framework to fine-tune VLM agents with RL and represents the SoTA method. Accordingly, we consider the method that solely clones the corrector’s thoughts, without RL (SFT-only). When evaluating final performance, we also include the LLaVA-mistral-7B base model and commercial API-based models. For Alfworld, the challenging environment with navigation and household tasks, we collect an expert dataset using a script policy and fine-tune several SoTA VLM models based on this dataset as baselines.
@@
The GTR framework leverages a corrector model with established capabilities for automating thought correction. In our experiment, we select the GPT-4o model for this role. Consistent with previous research, to ensure that RL training begins with a model possessing reasonable instruction following abilities, all training starts from an SFT-initialized LLaVA-mistral-7B model. We train the model for 15,000 steps on the Points24 task and 5,000 steps on the ALFWorld task, aligning with the baseline settings. On a single Nvidia (40GB) GPU, the training process with LoRA [14] takes approximately 30 hours to complete.